# AceMap: Knowledge Discovery through Academic Graph

### v1

@@ Abstract
The exponential growth of scientific literature requires effective management and extraction of valuable insights. While existing scientific search engines excel at delivering search results based on relational databases, they often neglect the analysis of collaborations between scientific entities and the evolution of ideas, as well as the in-depth analysis of content within scientific publications. The representation of heterogeneous graphs and the effective measurement, analysis, and mining of such graphs pose significant challenges. To address these challenges, we present AceMap, an academic system designed for knowledge discovery from a graph perspective. We present advanced database construction techniques to build the comprehensive AceMap database with large-scale academic publications that contain rich visual, textual, and numerical information. AceMap also employs innovative visualization, quantification, and analysis methods to explore associations and logical relationships among academic entities. AceMap introduces large-scale academic network visualization techniques centered on nebular graphs, providing a comprehensive view of academic networks from multiple perspectives. In addition, AceMap proposes a unified metric based on structural entropy to quantitatively measure the knowledge content of different academic entities. Moreover, AceMap provides advanced analysis capabilities, including tracing the evolution of academic ideas through citation relationships and concept co-occurrence, and generating concise summaries informed by this evolutionary process. In addition, AceMap uses machine reading methods to generate potential new ideas at the intersection of different fields. Exploring the integration of large language models and knowledge graphs is a promising direction for future research in idea evolution. Please visit https://www.acemap.info for further exploration.
@@ 2.1 AceMap Ontology
AceMap collects data from multiple disciplines, including more than 220 million papers, 103 million authors, 767 thousand research fields, 19 thousand academic institutes, 50 thousand journals, 4 thousand conferences, and 113 thousand affiliations. The AceMap Knowledge Graph (AceKG) provides comprehensive descriptions of 114.30 million academic entities, following a standardized ontology. In total, AceKG comprises a repository of 3.13 billion pieces of relationship information.
@@ 5 Quantification: Measure the amount of knowledge in papers
No.
Name
KQI (×$\times 10^{-4}$) No.
Name
KQI (×$\times 10^{-4}$)
1
Harvard University
4314.055 51
IBM
644.219
2
Stanford University
3073.220 52
Carnegie Mellon University
623.605
3
National Institutes of Health
2375.006 53
Purdue University
606.947
4
Massachusetts Institute of Technology
2207.661 54
McGill University
603.696
5
University of California, Berkeley
2193.094 55
Boston University
594.303
6
University of Michigan
1845.644 56
University of California, Irvine
591.930
7
University of Cambridge
1822.393 57
University of Edinburgh
586.073
8
Columbia University
1798.909 58
University of Manchester
586.070
9
University of California, Los Angeles
1782.025 59
University of Utah
580.203
10
University of Washington
1740.124 60
University of Melbourne
573.738
11
Yale University
1589.931 61
Osaka University
565.124
12
Max Planck Society
1563.456 62
Karolinska Institutet
552.533
13
University of Chicago
1499.324 63
Texas A&M University
548.657
14
University of Pennsylvania
1496.673 64
University of Rochester
540.651
15
University of Wisconsin-Madison
1409.335 65
University of Texas Southwestern Medical Center
537.137
16
Cornell University
1401.081 66
University of Iowa
537.066
17
University of California, San Diego
1373.094 67
Spanish National Research Council
526.036
18
University of Minnesota
1368.138 68
Michigan State University
523.051
19
Johns Hopkins University
1313.751 69
Case Western Reserve University
515.382
20
University of Oxford
1309.242 70
University of Virginia
512.322
21
Princeton University
1283.582 71
University of Colorado Boulder
506.716
22
New York University
1172.194 72
Vanderbilt University
506.472
23
University of California, San Francisco
1165.678 73
University of Copenhagen
500.675
24
University of Toronto
1161.220 74
Lund University
499.341
25
Duke University
1138.047 75
Brigham and Women’s Hospital
493.569
26
Centre national de la recherche scientifique
1115.222 76
Rockefeller University
473.213
27
University of Illinois at Urbana–Champaign
1078.021 77
University of Sydney
472.606
28
Northwestern University
1052.433 78
Brown University
467.860
29
University of Texas at Austin
1034.175 79
Arizona State University
463.401
30
Washington University in St. Louis
981.890 80
University of Amsterdam
460.407
31
University College London
971.396 81
King’s College London
456.631
32
University of Southern California
957.277 82
University of California, Santa Barbara
455.662
33
University of Tokyo
937.604 83
Hebrew University of Jerusalem
455.032
34
Chinese Academy of Sciences
903.581 84
Ludwig Maximilian University of Munich
448.234
35
California Institute of Technology
902.551 85
Utrecht University
447.436
36
University of North Carolina at Chapel Hill
882.683 86
University of Paris
446.511
37
University of Pittsburgh
853.370 87
Emory University
443.105
38
Pennsylvania State University
833.428 88
Katholieke Universiteit Leuven
439.741
39
Ohio State University
810.434 89
University of Bristol
435.826
40
University of California, Davis
798.666 90
University of Zurich
434.222
41
Rutgers University
759.650 91
University of Queensland
434.211
42
University of Florida
743.100 92
Australian National University
424.273
43
University of British Columbia
728.339 93
Baylor College of Medicine
420.760
44
Imperial College London
716.584 94
University of Alberta
419.484
45
French Institute of Health and Medical Research
712.008 95
University of Massachusetts Amherst
402.536
46
University of Maryland, College Park
701.468 96
Uppsala University
402.314
47
Mayo Clinic
666.846 97
Tohoku University
399.097
48
Bell Labs
664.646 98
National Institute of Standards and Technology
393.813
49
Kyoto University
662.686 99
University of Texas MD Anderson Cancer Center
393.462
50
University of Arizona
646.425 100
University of Miami
390.988
@@ 6.2 DeepReport: Discovering Innovative Concepts in Scientific Literature
Knowledge Conclusion by DeepReport Query: AlphaFold. Answer: AlphaFold is an artificial neural network that can be used to predict complex interactions between two or more heteromeric interfaces in real-time. AlphaFold, the deep learning algorithm developed by DeepMind, recently released the three-dimensional models of the whole human proteome to the scientific community. An AlphaFold model is trained specifically for multimeric inputs of known stoichiometry, which they call AlphaFold-Multimer, significantly increases accuracy of predicted multimeric interfaces over input-adapted single-chain AlphaFold while maintaining high intra-chain accuracy. These researches describe the ability of AlphaFold to predict intermolecular interactions before and after single mutations in proteins with respect to their predicted changes in amino acid binding sites. The algorithms are highly accurate and have enabled unprecedented expansion of the structural coverage of the known protein-sequence space.
@@ 7 Conclusion and Future Work
The AceMap system represents a step forward in academic data mining and knowledge discovery. We have extracted 19 million entities, 140 million relationships, and millions of figures, texts, and tables from the geoscience literature alone. Expanding our efforts to academic articles from other fields brings a massive scale and a wealth of unexplored knowledge waiting to be discovered. In addition, the dynamic nature of the scientific literature and the continuous evolution of AI technologies provide numerous opportunities for future research and improvements. Here, we outline some key directions for further exploration and development of AceMap.

### v2

@@ Abstract
The exponential growth of scientific literature requires effective management and extraction of valuable insights. While existing scientific search engines excel at delivering search results based on relational databases, they often neglect the analysis of collaborations between scientific entities and the evolution of ideas, as well as the in-depth analysis of content within scientific publications. The representation of heterogeneous graphs and the effective measurement, analysis, and mining of such graphs pose significant challenges. To address these challenges, we present AceMap, an academic system designed for knowledge discovery from a graph perspective. We present advanced database construction techniques to build the comprehensive AceMap database with large-scale academic entities that contain rich visual, textual, and numerical information. AceMap also employs innovative visualization, quantification, and analysis methods to explore associations and logical relationships among academic entities. AceMap introduces large-scale academic network visualization techniques centered on nebular graphs, providing a comprehensive view of academic networks from multiple perspectives. In addition, AceMap proposes a unified metric based on structural entropy to quantitatively measure the knowledge content of different academic entities. Moreover, AceMap provides advanced analysis capabilities, including tracing the evolution of academic ideas through citation relationships and concept co-occurrence, and generating concise summaries informed by this evolutionary process. In addition, AceMap uses machine reading methods to generate potential new ideas at the intersection of different fields. Exploring the integration of large language models and knowledge graphs is a promising direction for future research in idea evolution. Please visit https://www.acemap.info for further exploration.
@@ 2.1 AceMap Ontology
AceMap collects data from multiple disciplines, including more than 220 million papers metadata with 103 million authors, 767 thousand research fields, 19 thousand academic institutes, 50 thousand journals, 4 thousand conferences, and 113 thousand affiliations. The AceMap Knowledge Graph (AceKG) provides comprehensive descriptions of 114.30 million academic entities, following a standardized ontology. In total, AceKG comprises a repository of 3.13 billion pieces of relationship information.
@@ 5 Quantification: Measure the amount of knowledge in papers
No. Name KQI (×$\times 10^{-4}$) No. Name KQI (×$\times 10^{-4}$)
1 Harvard University 4314.055 51 IBM 644.219
2 Stanford University 3073.220 52 Carnegie Mellon University 623.605
3 National Institutes of Health 2375.006 53 Purdue University 606.947
4 Massachusetts Institute of Technology 2207.661 54 McGill University 603.696
5 University of California, Berkeley 2193.094 55 Boston University 594.303
6 University of Michigan 1845.644 56 University of California, Irvine 591.930
7 University of Cambridge 1822.393 57 University of Edinburgh 586.073
8 Columbia University 1798.909 58 University of Manchester 586.070
9 University of California, Los Angeles 1782.025 59 University of Utah 580.203
10 University of Washington 1740.124 60 University of Melbourne 573.738
11 Yale University 1589.931 61 Osaka University 565.124
12 Max Planck Society 1563.456 62 Karolinska Institutet 552.533
13 University of Chicago 1499.324 63 Texas A&M University 548.657
14 University of Pennsylvania 1496.673 64 University of Rochester 540.651
15 University of Wisconsin-Madison 1409.335 65 University of Texas Southwestern Medical Center 537.137
16 Cornell University 1401.081 66 University of Iowa 537.066
17 University of California, San Diego 1373.094 67 Spanish National Research Council 526.036
18 University of Minnesota 1368.138 68 Michigan State University 523.051
19 Johns Hopkins University 1313.751 69 Case Western Reserve University 515.382
20 University of Oxford 1309.242 70 University of Virginia 512.322
21 Princeton University 1283.582 71 University of Colorado Boulder 506.716
22 New York University 1172.194 72 Vanderbilt University 506.472
23 University of California, San Francisco 1165.678 73 University of Copenhagen 500.675
24 University of Toronto 1161.220 74 Lund University 499.341
25 Duke University 1138.047 75 Brigham and Women’s Hospital 493.569
26 Centre national de la recherche scientifique 1115.222 76 Rockefeller University 473.213
27 University of Illinois at Urbana–Champaign 1078.021 77 University of Sydney 472.606
28 Northwestern University 1052.433 78 Brown University 467.860
29 University of Texas at Austin 1034.175 79 Arizona State University 463.401
30 Washington University in St. Louis 981.890 80 University of Amsterdam 460.407
31 University College London 971.396 81 King’s College London 456.631
32 University of Southern California 957.277 82 University of California, Santa Barbara 455.662
33 University of Tokyo 937.604 83 Hebrew University of Jerusalem 455.032
34 Chinese Academy of Sciences 903.581 84 Ludwig Maximilian University of Munich 448.234
35 California Institute of Technology 902.551 85 Utrecht University 447.436
36 University of North Carolina at Chapel Hill 882.683 86 University of Paris 446.511
37 University of Pittsburgh 853.370 87 Emory University 443.105
38 Pennsylvania State University 833.428 88 Katholieke Universiteit Leuven 439.741
39 Ohio State University 810.434 89 University of Bristol 435.826
40 University of California, Davis 798.666 90 University of Zurich 434.222
41 Rutgers University 759.650 91 University of Queensland 434.211
42 University of Florida 743.100 92 Australian National University 424.273
43 University of British Columbia 728.339 93 Baylor College of Medicine 420.760
44 Imperial College London 716.584 94 University of Alberta 419.484
45 French Institute of Health and Medical Research 712.008 95 University of Massachusetts Amherst 402.536
46 University of Maryland, College Park 701.468 96 Uppsala University 402.314
47 Mayo Clinic 666.846 97 Tohoku University 399.097
48 Bell Labs 664.646 98 National Institute of Standards and Technology 393.813
49 Kyoto University 662.686 99 University of Texas MD Anderson Cancer Center 393.462
50 University of Arizona 646.425 100 University of Miami 390.988
@@ 6.2 DeepReport: Discovering Innovative Concepts in Scientific Literature
Knowledge Conclusion by DeepReport
Query: AlphaFold. Answer: AlphaFold is an artificial neural network that can be used to predict complex interactions between two or more heteromeric interfaces in real-time. AlphaFold, the deep learning algorithm developed by DeepMind, recently released the three-dimensional models of the whole human proteome to the scientific community. An AlphaFold model is trained specifically for multimeric inputs of known stoichiometry, which they call AlphaFold-Multimer, significantly increases accuracy of predicted multimeric interfaces over input-adapted single-chain AlphaFold while maintaining high intra-chain accuracy. These researches describe the ability of AlphaFold to predict intermolecular interactions before and after single mutations in proteins with respect to their predicted changes in amino acid binding sites. The algorithms are highly accurate and have enabled unprecedented expansion of the structural coverage of the known protein-sequence space.
@@ 7 Conclusion and Future Work
The AceMap system represents a step forward in academic data mining and knowledge discovery. We have extracted large-scale entities, relationships, and figures, texts and tables from Open Access (OA) papers with the collaboration with geoscientists. Expanding our efforts to academic articles from other fields brings a massive scale and a wealth of unexplored knowledge waiting to be discovered. In addition, the dynamic nature of the scientific literature and the continuous evolution of AI technologies provide numerous opportunities for future research and improvements. Here, we outline some key directions for further exploration and development of AceMap.